Company mass produces chairs and has many colors but only one design. Design is the company's key to success and there seems to be no end for demand

The following data were extracted from the company's standard cost sheet

Plastic: 10 pounds at $4.50 per pound

Molding: 3 feet at $3.00 per foot

Direct labor: 4 hours at $6.00 per hour

Variable Overhead: $3 per direct labor hour

Fixed Overhead: $55,000 per period

Transactions during the month of June were as follows

Purchased plastic at $4.45 per pound and issued 185,000 pounds to production

Purchased molding at $3.10 per foot and issued 50,000 feet to production

The direct labor payroll totaled $435,000 for 72,000 hours

Total overhead costs were $275,000, including $221,125 of variable overhead

Produced 18,000 chairs during the month

Required:

Calculate all material, labor, variable overhead, and fixed overhead variances

Interpret the material and labor variances, what do they indicate about the company's performance?

How could the company control or better manage its operation

In your opinion what are the best options and why?   
Answer #1
1] Direct material price variance = Actual quantity*(Actual price-Standard price)
For Plastic = 185000*(4.45-4.5) = $            9,250 Favorable
For Molding = 50000*(3.1-3) = $            5,000 Unfavorable
Direct material price variance $            4,250 Favorable
[As there are two materials, the variances would be the
sum of the individual variances]
Direct material quantity variance = Standard price*(Actual quanity-Standard quantity)
For Plastic = 4.5*(185000-18000*10) = $          22,500 Unfavorable
For Molding = 3*(50000-18000*3) = $          12,000 Favorable
Direct material quantity variance $          10,500 Unfavorable
Direct material cost variance = Direct material price variance+Direct material quantity variance = 10500 [U]+4250[F] = $            6,250 Unfavorable
Also, (185000*4.45+50000*3.1)-(18000*10*4.5+18000*3*3) = $            6,250 Unfavorable
2] Direct labor rate variance = Actual hours*(Actual rate-Standard rate) = 72000*(435000/72000-6) = $            3,000 Unfavorable
Direct labor efficiency variance = Standard rate*(Actual hours-Standard hours) = 6*(72000-18000*4) = 0 NA
Direct labor cost variance = 3000 [U]- 0 = $            3,000 Unfavorable
Also, 435000-18000*4*6 = $            3,000 Unfavorable
3] VOH rate variance = Actual hours*(Actual rate-Standard rate) = 72000*(221125/72000-3) = $            5,125 Unfavorable
VOH efficiency variance = Standard rate*(Actual hours-Standard hours) = 3*(72000-18000*4) = 0 NA
VOH cost variance = 5125 [F]- 0 = $            5,125 Unfavorable
Also, 221125-18000*4*3 = $            5,125 Unfavorable
4] Fixed overhead spending variance = Actual fixed overhead-Budgeted overhead = (275000-221125)-55000 = $            1,125 Favorable
Fixed overhead volume variance = Budgeted overhead-Overhead assigned = Cannot be calculated as the planned production is not available to ascertain the predetermined overhead rate.
5] The direct material price variance is favorable, but, the
quantity variance is unfavorable. It may be due to low
quality material that was purchased, which, has resulted in
favorable price variance and unfavorable quantity variance.
If the material were good, the unfavorable quantity variance
should be investigated to pin point inefficiency in processing.
It may also be due to poor workmanship.
The direct labor rate variance is unfavorable which may be
due to increase in wage rates. The firm should try to
negotiate better or increase the efficiency of the labor.
The VOH rate variance is unfavorable. The expenditure
should be analyzed to reduce the incurrence.
Fixed overhead spending should also be analyzed item wise.